### v3

# Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning

Abstract
Humans perceive the world as a series of sequential events, which can be hierarchically organized with different levels of abstraction based on conceptual knowledge. Drawing inspiration from this human cognition aspect, we explore how human tendencies to organize and connect information can enhance the training of deep learning models. Specifically, we propose a novel approach to mitigate catastrophic forgetting by exploiting the relationships between continuously emerging class data. By building a hierarchical tree structure based on the expanding set of labels, we gain fresh insights into the data, identifying groups of similar classes could easily cause confusion. Additionally, we delve deeper into the hidden connections between classes by exploring the original pretrained model behavior through an optimal transport-based approach. From these insights, we propose a novel regularization loss function that encourages models to focus more on challenging knowledge areas, thereby enhancing overall performance. Experimentally, our method demonstrated significant superiority over state-of-the-art methods on various continual learning benchmarks. Our code is available at https://anonymous.4open.science/r/HierC-089B/.
*Hoang Phan and Minh Le contributed equally to this work
1 Introduction
Continual Learning (CL) [45, 27] is a research direction that focuses on realizing the human dream of creating truly intelligent systems, where machines can learn on the fly, accumulate knowledge, and operate in constantly changing environments as a human’s companion. Despite the impressive capabilities of A.I systems, continual learning remains a challenging scenario due to the tendency to forget obtained knowledge when facing new ones, known as catastrophic forgetting [12]. In dealing with this challenge, traditional CL methods often rely on storing past data for replaying during new tasks [28, 6], which can raise concerns about memory usage and privacy. Besides, prior work shows that replay methods result in overfitting and poor generalization [28, 42, 7]. To overcome these limitations, recent methods proposed leveraging the strong generalization ability of pretrained models [14, 19] to solve sequences of CL tasks. A notable line of work is the prompt-based approach [48, 38, 25], where a small set of learnable prompts is injected to pretrained backbones for adapting emerging tasks over time.
While these prompt-based methods have demonstrably achieved impressive results, they only consider forgetting caused by changes of common parameters across tasks during learning [49, 48] or the inherent mismatch between the chosen prompts at training and testing [44, 40, 53, 38]. In this work, we further complement these views by showing that the forgetting of old tasks can potentially come from the uncontrolled overlap between old and new emerging class representations in the latent space. That is, models become more confused in distinguishing these classes, resulting in performance degradation w.r.t previous tasks over time (i.e., forgetting). Furthermore, we find that existing methods only utilize limited information from the training dataset and often treat classes equally during training. Consequently, they overlook the opportunities to enhance the distinguishability of models, especially between the class representations of old and new tasks, and thus, hinder the models’ ability to mitigate forgetting [11, 21].
In addition, we discover that human natural learning behavior has many valuable aspects, especially the habit of analyzing data, organizing them in a meaningful way, and finding connections between old and new knowledge [37, 3, 39, 30]. Inspired by these practices, we investigate the characteristics of current common benchmark datasets, as well as the behavior of pretrained models. Our findings reveal that the data can always be categorized into consistent groups, regardless of their arrival times. Each group usually includes classes with similar semantic characteristics that models may confuse and, thus, should be paid special attention to during training. Identifying these kinds of groups will benefit in improving model representation learning, as well as aiding in addressing catastrophic forgetting.
Therefore, we propose a training strategy that constantly considers emerging data in groups, following a hierarchical tree-like taxonomy developed based on expert knowledge. In particular, during the training on new tasks, the feature extractor is trained to distinguish all classes so far, especially focusing on classes within the same group. We observe that images belonging to concepts/labels within the same leaf group in the taxonomy share strong visual and semantic correlations, leading to overlapping representations in the latent space, which compromises performance. By encouraging the feature extractor to separate and contrast their representations more distinctly, we can reduce the overlap of easily confused classes, thereby further improving performance. This strategy not only mitigates forgetting when new classes emerge but also consolidates domain-specific knowledge in each of these leaf groups.
Contribution. We introduce a method named Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning (TCL). Our main contributions are summarized as follows:
  • We conduct empirical analysis to suggest a new cause of catastrophic forgetting in prompt-based continual learning models, which inherently comes from the uncontrolled growth of incoming classes on the latent space.
  • Originating from the research findings of cognitive science, we propose a novel approach to reduce forgetting by exploiting relationships between data. By dynamically constructing label-based hierarchical taxonomies and leveraging the prior knowledge of pretrained models via an optimal transport-based approach, we can identify the challenging knowledge areas that need further focus during the sequence of tasks.
  • We empirically evaluate the effectiveness of our method against current state-of-the-art pretrained-based baselines across various continual learning benchmarks.
Organization. The rest of the paper is structured as follows. In Section 2, we present related work. Then in Section 3, we formulate the problem and introduce a new perspective to explain the cause of forgetting in prompt-based CL models. Section 4 transitions from the motivation provided by Cognitive Science insights to the proposed training strategy, emphasizing the importance of leveraging relationships between class data. Section 5 presents the experimental results, and finally, we discuss the limitations and suggest future directions in Section 6.
2 Related Work
Class Incremental Learning.
This is one of the most challenging and widely studied CL scenarios [41, 44], where task identity is unknown during testing, and data of previous data is inaccessible during current training [29, 36, 16, 13]. This work follows the setting of CIL and proposes a novel approach to mitigate forgetting and improve performance for prompt-based CL models.
Prompt-based Continual Learning.
This line of work exploits the power of pretrained backbone to quickly adapt to the sequence of downstream tasks by updating just a small number of parameters or prompts. Initial work like [49, 48, 38] typically assign a set of prompts to tasks, enhancing the adaptability of the backbone to downstream tasks. However, the absence of explicit constraints can lead to feature overlapping between classes from different tasks. Therefore, recent methods employ some types of contrastive loss [44, 24] or utilize Vision Language models [46, 33] to better separate features from tasks. However, they treat all classes equally during training, missing the opportunity to learn in challenging areas where classes have many similarities and are easily confused. In this work, we propose a novel approach to exploit the relationships within data, allowing the model to recognize groups of these classes, and develop a deeper understanding of the respective knowledge areas, thereby reducing forgetting and enhancing its ability to learn new tasks.
3 Preliminary Analysis
In this section, we first present the problem formulation of prompt-based continual learning in Section 3.1. Following this, we analyze the sources of forgetting in these methods in Section 3.2.
3.1 Problem formulation
We consider the Class Incremental Learning setting [54, 28, 44], where a model has to learn from a sequence of $T$ visual classification tasks without revisiting old task data during training or accessing task IDs during inference. Each task $t\in\{1,...,T\}$ has a respective dataset $\mathcal{D}_{t}$, containing $n_{t}$ i.i.d. samples ${(\bm{x}_{t}^{i},y_{t}^{i})}_{i=1}^{n_{t}}$.
In this work, we design our model as a composition of two components: a pretrained Vision Transformer (ViT) [9] backbone $f_{\Phi}$ and a classification head $h_{\psi}$. That is, we have the model parameters $\theta=(\Phi,\psi)$. Similar to other existing prompt-based methods [44, 38], we incorporate into the pretrained ViT backbone a set of prompts $\bm{P}$. We denote the overall network after incorporating the prompts as $f_{\Phi,\bm{P}}$.
Dataset Sup-21K DINO-1K
Split-Imagenet-R 3.82 80.31
Split-CIFAR100 1.32 48.43
Split-CUB-200 2.04 63.56
(a) Distribution shift of representations corresponding to $\mathcal{D}_{1}$ after learning sequence of tasks.
(b) Within task accuracy on $\mathcal{D}_{1}$ (Sup-21K)
(c) True accuracy on $\mathcal{D}_{1}$ (Sup-21K)
Figure 1: Empirical study about forgetting of HiDe-Prompt. Table (a) reports L2 Wasserstein distance between the latent distributions corresponding to $\bm{P}_{t}$ and $\bm{P}_{\hat{t}}$ of data task 1, after learning the final task; Figure (b) Accuracy of the first task over time, using classification head $s(\bm{x})$; Figure (c) Accuracy of the first task over time, using classification head $h_{\psi}$.
3.2 Forgetting in Prompt-based Methods
In the context of Continual Learning, forgetting refers to the phenomenon in which performance on previously learned tasks decreases over time. Current prompt-based CL methods, which leverage the power of pretrained models, attribute forgetting either to (I) changes in parameters from the backbone when using the common prompt pool $\bm{P}$ for all tasks [49, 48] or to (II) the inherent mismatch between the models used at training and testing, as discussed and analyzed in [53, 40]. In this work, we complement these views by providing empirical studies to offer new insights about the reason for forgetting in this type of model, which arises from overlapping between old and new emerging class representations.
To uncover the new factor causing forgetting, we propose to consider the cases where the behavior of backbone w.r.t each learned task is nearly unchanged over time - meaning (I) and (II) roughly would not happen. Firstly, to eliminate concerns about changing learned parameters (I), we consider methods that propose using a distinct set of prompts $\bm{P}_{t}$ to a specific task $t$. Then, the remaining potential factor is the difference between the prompt chosen at inference time, which we refer as $\bm{P}_{\hat{t}}$, and the one used during training $\bm{P}_{t}$. Thus, we conduct experiments on HiDe-Prompt [44], which is the latest SOTA in prompt-based CL, to measure the differences between the features formed when using these two prompts. Particularly, we consider $\bm{x}\in\mathcal{D}_{1}$ that belongs to the first task and measure the L2 Wasserstein distance $W_{2}(Q,\hat{Q})$ [20] between $Q$ (i.e., the latent distribution corresponding to $\bm{P}_{t}$, which consists of $f_{\Phi,\bm{P}_{{t}}}(\bm{x})$ for $\bm{x}\sim\mathcal{D}_{1}$) and $\hat{Q}$ (i.e., the latent distribution corresponding $\bm{P}_{\hat{t}}$ after learning the last task, consists of $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$ for for $\bm{x}\sim\mathcal{D}_{1}$). The results in Table 1(a) show that the difference between these distributions is negligible when using Sup-21K as a pre-trained backbone, in stark contrast to the significant shift observed in DINO-1K.
For a detailed examination of Sup-21K, we propose an experimental setup to analyze the performance degradation of corresponding models over time. Specifically, we train a classifier $s(\bm{x})$ on the latent space $f_{\Phi,\bm{P}_{1}}(\bm{x})$ derived from the first task $\mathcal{D}_{1}$ to accurately classify the classes within this initial task. As the model undergoes continual training on a sequence of subsequent tasks, we assess the ability of $s(\bm{x})$ to classify instances from $\mathcal{D}_{1}$ based on the latent space representation $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$, defining this as within-task accuracy. This is compared to the true accuracy when using the classifier $h_{\psi}$. The results in Figure 1(b), show that within-task accuracy remains nearly stable, suggesting minimal shift between representations $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$ and $f_{\Phi,\bm{P}_{t}}(\bm{x})$. In contrast, Figure 1(c), illustrates a significant decrease in true accuracy, raising the question of whether we have overlooked additional factors contributing to final forgetfulness.
Additionally, considering the inference feature space of $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$, we can see that as more tasks arrive, the number of classes increase, making the space fuller and increasing the possibility of overlap between class distributions. To support this point, we illustrate the representation of some class images in Figure 2, as well as the corresponding t-SNE visualizations in Figure 4. In particular, after Task 1, we have representations of ”oak tree”, ”mouse” and ”porcupine” located in quite separate locations. However, when Task 2 and then Task 3 arrive, the appearance of ”willow tree” and ”pine tree” makes the latent space become fuller, and ”oak tree” no longer maintains the separation from the remaining classes as before, leading to a remarkable drop in model performance.
Based on the evidence above, another key cause of catastrophic forgetting in prompt-based continual learning that should be recognized is the addition of new classes, which gradually fills the latent space and may overlap with existing ones. The overlap causes confusion in distinguishing between old and new classes, thereby reducing the performance of learned tasks over time. This is coincidently similar to the behavior of human memory, where although the old knowledge exists somewhere in the brain, still be confused due to the newly acquired information [1, 50, 26, 31, 51]. While existing CL methods emphasize the importance of representation learning to keep classes distinct [44, 24], none explicitly acknowledge the overlap between new and old tasks as a source of forgetting. Recognizing this motivates us to propose a novel method, focusing on identifying easily confused class pairs, thereby reducing forgetting and improving performance.
Figure 2: Problem: When new classes arrive, the latent space of a model for all tasks so far becomes fuller, and class representations tend to be overlapped, leading to performance degradation in old tasks. Our solution: In this work, we focus on separating easily confused classes whose concepts/labels lie in the same leaf group on a hierarchical taxonomy, suggested by expert knowledge (i.e., ChatGPT - Figure 3)
4 Proposed Method
In the previous section, we noted that increasing overlap in data representations as more tasks arrive is one of the main reasons leading to performance degradation (i.e., forgetting). Therefore, if we can identify easily confused classes/concepts, we will better enhance the distinguishability of models and thus reduce forgetting. Inspired by Cognitive Science studies, regarding the benefits of organizing information in a meaningful way (Section 4.1), we propose using expert knowledge to structure the concepts/labels of CL tasks in the corresponding hierarchical taxonomies. Interestingly, we find that concepts within the same group in each taxonomy tend to be visually and semantically similar, potentially causing more overlap in the latent space and confusion for the CL classifier. Motivated by this observation, we propose group-based contrastive learning in Section 4.2 to maximize the separability of these concepts, thereby mitigating forgetting over time. Additionally, to reinforce the effectiveness of this strategy, Section 4.3 presents an optimal transport-based strategy, which harnesses prior knowledge of pretrained backbones from a new perspective.
4.1 Motivation
Insights from Cognitive Science.
Research in Cognitive Science highlights the importance of reflection, and organization of information as critical components for effective learning. Studies show that when learners take time to reflect on their experiences, they deepen their understanding and enhance retention [37]. This reflective practice encourages individuals to connect new information with existing knowledge, fostering a more integrated learning experience [3]. Linking new material with relevant prior knowledge—often referred to as associative learning—further strengthens memory retention [39] and also benefits future learning. Moreover, organizing information into coherent structures, such as outlines or concept maps, allows learners to see relationships between concepts, making it easier to retrieve the information later [30].
Figure 3: The hierarchical taxonomy obtained when learning Task 3 on Split-CIFAR100.
(a) Training tasks independently with $\mathcal{L}_{CE}$
(b) Ours (TCL)
Figure 4: t-SNE visualization of classes within leaf groups of Four-legged animals ($\bullet$ circular points) and Plants ($\blacktriangle$ triangular points) when learning Task 3, Split-CIFAR100. The order of appearance of the classes: Task 1 - ”mouse”, ”porcupine”, ”oak tree”; Task 2 - ”willow tree”; Task 3 - ”otter”, ”hamster”, ”pine tree” (Figure 2 and 3). We can see that if we train tasks independently with $\mathcal{L}_{CE}$, the classes within each such group, which arrive at different tasks, can be overlapped seriously (Figure 4(a)).
Our Approach.
It is evident that besides reflection - comparison of old and new knowledge, the key factor in learning efficiently is to organize and link information in an insightful way, where concepts are linked and arranged according to their semantic meanings. This observation motivates us to develop deep learning models that learn concepts structured in a hierarchical taxonomy. The aim is to enable the models to grasp relevant concepts more effectively, thereby enhancing distinguishability and mitigating catastrophic forgetting. More specifically, we propose structuring the data labels in a tree-like taxonomy, which can be flexibly expanded over time using domain expertise. Based on this structure, we have a view of the relationships between classes, identifying which classes belong to the group with many shared characteristics, easily confused, and require more focus (see Figures 3, 4).
Taking the learning process of Split-CIFAR100 (see Appendix A.1) as an example, when training on task $t=3$, we can construct a tree-like taxonomy of concepts/labels, as shown in Figure 3. We observe that the classes under a same leaf in the structure (e.g., oak tree, willow tree, and pine tree in the leaf group ”plants”) exhibit stronger visual and semantic correlations. Consequently, as shown in Figure 4a, these class features are more overlapping with each other rather than with another class from other leaf groups (e.g., ”four-legged”). In this way, the tree-like taxonomy serves as a tool to help identify easily confused classes, facilitating the subsequent process of making them more separable in the feature space. Leveraging the insight obtained from the taxonomy above, we aim to train a network so that all class representations must be distinct, especially those within each leaf group. Over time, this strategy not only helps the model better mitigate forgetting caused by representation overlap but also reinforces existing knowledge to facilitate future learning.
4.2 Leveraging Hierarchical Taxonomies for Better Representation Learning
During the training process, whenever a new class appears, its label name is automatically added to the tree-like taxonomy, into a leaf group containing classes with similar characteristics (Figure 3). To develop this hierarchical structure, we can rely on expert knowledge, for example, ChatGPT, which can help us incrementally construct a meaningful and semantically related tree (see Appendix B). Structuring information in this way not only aligns with how the human brain effectively connects and remembers information but also provides useful insights during training, indicating how each knowledge is related to the other and which requires further focus.
For each task $t$, we dedicate a set of prompt $\bm{P}_{t}$. In each task, we aim to minimize overlap between all classes so far, especially focusing on increasing the separability between classes belonging to the same leaf group on the taxonomy (e.g., four-legged mammals, plants, etc.,). That is, taxonomy acts as a reference information channel to support the training process. Let $g$ be a leaf group on the taxonomy $\mathcal{G}$ (i.e., $g\in\mathcal{G}$), ${X}^{g}_{k}$ and ${Y}^{g}_{k}$ denote the corresponding sets of input samples and labels under the group $g$ that belong to the task $k$ ($k\leq t$). Besides Cross Entropy loss $\mathcal{L}_{CE}$, we propose using a regularization loss function for sample ${\bm{x}}$ that arrives in task $t$ and belongs to leaf group $g$ as follows:
$\mathcal{L}_{\mathcal{G}}(\psi,\bm{P}_{t},\bm{x})=\alpha\mathcal{L}_{g}(\psi,\bm{P}_{t},\bm{x})+\beta\mathcal{L}_{all}(\psi,\bm{P}_{t},\bm{x})$ (1)
where we have defined
$\mathcal{L}_{g}(\psi,\bm{P}_{t},\bm{x})$ $=-\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X^{g}_{\overline{1,t}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$
$\displaystyle\mathcal{L}_{all}(\psi,\bm{P}_{t},\bm{x})$ $=-\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X_{\overline{1,t}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$
are the Supervised Contrastive loss that we put on class representation within leaf group $g$, and all classes so far, respectively; $u({\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}^{\prime}})=\exp(\frac{{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}^{\prime}}}{\tau})$, with ${\bm{z}}_{\bm{x}}=f_{\Phi,\bm{P}_{t}}({\bm{x}})$ is the feature vector on the latent space of the prompt-based model and $y_{\bm{x}}$ is the ground truth label of ${\bm{x}}$, $\tau$ is the temperature with $\tau=0.1$ for all experiments, and $\alpha$ is the coefficient that controls how much we want to force classes within leaf group $g$ stay apart further.
In addition, because the data of old tasks can not be accessed when training the new one, we encode the information of each trained class $c$ in the form of a Gaussian Mixture model $\text{GMM}_{c}=\{\mathcal{N}(\mu_{c,i},\Sigma_{c,i})\}_{i=1}^{K}$, at the end of the task, where $K$ is the number of components. Then in task $t$, the representation of each old class $c$ is sampled from $\text{GMM}_{c}$, to compute loss functions in Eq. (1), supporting the training of the current task.
Looking closer at Eq. (1), this equation implies that when learning a new task, the model will be encouraged to identify the decision boundary between old and new classes, especially focusing on those in the same leaf group (see Figure 2). This manipulation matches the way our brain naturally works - without reflection and comparison, we would be confused about old and new concepts, especially those sharing many common characteristics, leading to incorrect judgments and decisions in practice. For A.I models, learning new tasks without thoroughly considering the learned information of old tasks can lead to uncontrolled overlapping of layer representations in the latent space, resulting in forgetting and harming final performance. In this work, thanks to the insight from the hierarchical taxonomy, we can identify which group of classes that easy to get overlapped in the latent space, thereby actively intervening and enhancing the model representation learning.
Figure 5: Wasserterin distance between classes (Split-CIFAR100) in latent space of pretrained backbone (Sup-21K).
4.3 Harnessing Prior Knowledge from Pretrained Models via Optimal Transport
Considering a leaf group, there may be data classes with varying levels of overlap in the latent space. For example, Figure 5 provides statistical results regarding the relative position of classes in leaf groups ”plants” and ”four-legged mammals”. We can see that in the group of animals, the L2-Wasserstein distance between class representations of ”mouse” and ”otter” (39.30) is more significant than the one between ”mouse” and ”hamster” (26.05), meaning that the class representation of the second pair can overlap more than the first one. Although focusing on classes within the same leaf group helps improve the ability to recognize difficult-to-identify classes, we still treat all classes in that group equally. Thus, the algorithm may inadvertently ignore the pairs of classes that are easily confused and need to be further distinguished.
Moreover, pretrained models are extensively trained on large datasets, which results in their substantial generalization abilities. That is the reason why they are considered as good starting points for the adaptation to downstream tasks [8, 4, 17, 15, 35]. In this work, we propose to exploit the pretrained models from a novel perspective, which can comprehend the use of the label-based tree-like taxonomy during training. In particular, we take advantage of these models to obtain prior assumptions about the relationships between the class representations on the latent spaces, thereby elegantly intervening in better separating each pair of classes within a leaf group.
Method Split CIFAR-100 Split ImageNet-R 5-Datasets Split CUB-200
FAA ($\uparrow$) FFM ($\downarrow$) FAA ($\uparrow$) FFM ($\downarrow$) FAA ($\uparrow$) FFM ($\downarrow$) FAA ($\uparrow$) FFM ($\downarrow$)
L2P 83.06 ±$\pm 0.17$ 6.58 ±$\pm 0.40$ 63.65 ±$\pm 0.12$ 7.51 ±$\pm 0.17$ 81.84 ±$\pm 0.95$ 4.58 ±$\pm 0.53$ 74.52 ±$\pm 0.92$ 11.25 ±$\pm 0.23$
DualPrompt 86.60 ±$\pm 0.19$ 4.45 ±$\pm 0.16$ 68.79 ±$\pm 0.31$ 4.49 ±$\pm 0.14$ 77.91±$\pm 0.45$ 13.17 ±$\pm 0.71$ 82.05±$\pm 0.95$ 3.56 ±$\pm 0.53$
OVOR 86.68 ±$\pm 0.22$ 5.25 ±$\pm 0.12$ 75.61 ±$\pm 0.82$ 5.77 ±$\pm 0.12$ 82.34 ±$\pm 0.48$ 4.83 ±$\pm 0.35$ 78.12 ±$\pm 0.0.65$ 8.13 ±$\pm 0.52$
S-Prompt++ 88.81 ±$\pm 0.18$ 3.87 ±$\pm 0.05$ 69.68 ±$\pm 0.12$ 3.29 ±$\pm 0.05$ 86.19±$\pm 0.65$ 4.67 ±$\pm 0.72$ 83.12 ±$\pm 0.54$ 2.72 ±$\pm 0.64$
CODA-Prompt 86.94 ±$\pm 0.63$ 4.04 ±$\pm 0.18$ 70.03 ±$\pm 0.47$ 5.17 ±$\pm 0.22$ 64.20 ±$\pm 0.53$ 17.22 ±$\pm 0.55$ 74.34 ±$\pm 0.68$ 12.05 ±$\pm 0.41$
CPP 91.12 ±$\pm 0.12$ 3.33 ±$\pm 0.18$ 74.88 ±$\pm 0.07$ 4.08 ±$\pm 0.03$ 92.92 ±$\pm 0.17$ 0.23 ±$\pm 0.07$ 82.35 ±$\pm 0.23$ 3.24 ±$\pm 0.32$
HiDe-Prompt 92.61 ±$\pm 0.28$ 1.52 ±$\pm 0.10$ 75.06 ±$\pm 0.12$ 4.05 ±$\pm 0.19$ 93.92 ±$\pm 0.33$ 0.31 ±$\pm 0.12$ 86.62 ±$\pm 0.35$ 2.55 ±$\pm 0.15$
Ours (TCL) 93.94 ±$\pm 0.23$ 1.05 ±$\pm 0.15$ 76.41 ±$\pm 0.12$ 4.02 ±$\pm 0.22$ 95.02 ±$\pm 0.20$ 0.21 ±$\pm 0.15$ 87.93 ±$\pm 0.22$ 2.01 ±$\pm 0.23$
Table 1: Overall performance comparison. We provide FAA and FFM of all methods, with standard deviation taken over at least 3 runs of different random seeds. The results corresponding to the best FAA among baselines are underlined.
Firstly, to extract the relationship between the classes, we compute the L2-Wasserstein distance of feature distributions of each class pair. Specifically, we denote $D_{c_{i}}^{\Phi}$ be the distribution of class $c_{i}$ on the latent space of the pretrained model $f_{\Phi}$ to distinguish it from the class distribution on prompted latent space $f_{\Phi,\bm{P}}$, as described in Section 4.2. The distribution $D_{c_{i}}^{\Phi}$ is obtained in the form of a Gaussian Mixture model at the end of the respective task, only once and then kept fixed. In this way, when $t$ tasks have arrived, we have the corresponding sets of distributions $\{D_{c}^{\Phi}\}_{c\in Y_{1,t}}$ of all $m_{t}$ classes so far. Therefore, we gradually complete the WD-based matrix between pairs of classes:
$M=[W_{2}(D_{c_{i}}^{\Phi},D_{c_{j}}^{\Phi})]_{m_{t}\times m_{t}}.$ (2)
The fact is that the smaller the $W_{2}$ between the representations of two classes, the closer they are, and the more force can be needed to separate them. Therefore, we employ these results to obtain the corresponding weight factors, which control how much constraint needs to be applied to each pair. In particular, we compute the weight matrix:
$\Gamma=\left[\gamma_{ij}\right]_{m_{t}\times m_{t}}=\left[1/{\exp(M_{ij}/\delta)}\right]_{m_{t}\times m_{t}}$ (3)
where $\delta$ is a temperature. We then apply this information to obtain a weighted version of $\mathcal{L}_{g}$, in which the closer the two class distributions are, the larger the weight assigned, and they will be focused to push away. Consequently, our regularization loss becomes:
$\mathcal{L}_{g}(\cdot)=-\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X^{g}_{\overline{1,t}}}\gamma_{y_{x}y_{\tilde{x}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$ (4)
This strategy is completely economical and aligns well with the CL learning scheme, as the distance between representations of each class pair only needs to be computed once, and the matrix $M$ is continuously expanded when new classes arrive. Practically, when learning a new task, the first epoch is for capturing information about the behavior of the pretrained model on the data of this task. Moreover, this approach is also aligned with the findings in Cognitive Science [34, 2], showing that the accumulated experiences from past learning create momentum for learning new skills more effectively.
Finally, our objective function during training can be formulated as follows:
$\mathcal{L}=\mathcal{L}_{CE}+\alpha\mathcal{L}_{g}+\beta\mathcal{L}_{all}$ (5)
5 Experiments
5.1 Experimental setup
Datasets. We use 4 common CIL benchmarks, including Split CIFAR-100, Split ImageNet-R, 5-Datasets, and Split CUB-200 (see Appendix A.1).
Baselines. We compare ours with the latest state-of-the-art prompt-based methods, including the methods using shared prompts for all tasks: L2P [49], DualPrompt [48], OVOR [18]; and the methods dedicates a distinct prompt set for each task like: S-Prompt++ [47], CODA-Prompt [38], HiDe-Prompt [44], CPP [25] (see Appendix A.2).
Metrics. We use two main metrics, including the Final Average Accuracy (FAA) and the Final Forgetting Measure (FFM), denoting the average accuracy and forgetting of all tasks after learning the sequence of tasks, respectively (see Appendix A.2). Besides, the detailed implementation is described in Appendix A.4.
(a) The role of $\mathcal{L_{G}}$
(b) The role of OT-based strategy
(c) Varying the value of $\alpha,\beta$
Figure 6: Ablation studies about our training strategy.
5.2 Experimental result
Our approach achieves superior results compared to baselines.
Table 1 presents the overall performance comparison between our proposed method and all the baselines. The key observation is that our method is the strongest one, by the gap between our method and the runner-up method being around 1.5% in terms of FAA on all considered datasets. Additionally, the results show that our method avoids forgetting better than all baselines, by the gap up to 17% on 5-Datasets.
Ablation studies about our training strategy
Figure 6 reports the ablation studies demonstrating the effectiveness of our training strategy. Particularly, compared to training tasks independently using Cross Entropy loss $\mathcal{L}_{CE}$ like in DualP, L2P, and CODA-P, exploiting the relationships between data classes with the label-based hierarchical taxonomy and the WD-based cost matrix helps improve FAA by about 5% to 10% (Figure 6(a)). Besides, when examining the role of exploiting additional prior information from pretrained backbones using the OT approach, we see that FAA is improved from 0.6% to 0.8% (Figure 6(b)). These results demonstrate the positive impact of this component, confirming the importance of exploiting correlations between class data during training. In both figures, the improvements on Split-CIFAR100 and 5-Datasets are the lowest, while it is more pronounced on Split-CUB-200. This may be because the groups of these two datasets (Split-CIFAR100 and 5-Datasets) have fewer overlapping classes, as the classes in each group likely have more recognizable features. Meanwhile, Split-CUB-200 is a dataset of birds with images that can be difficult for human eyes to recognize, thus so our method performs better.
In addition, Figure 6(c) provides the experimental results on the dataset Split-CUB-200, when varying $\alpha$ and $\beta$, which control the intensification of impact on each leaf group of $\mathcal{L_{G}}$ during training. The data shows that with a large enough value of $\beta$, our TCL is not sensitive to $\alpha$ within its acceptable range. Conversely, if $\alpha$ is small, the quality of the model can change more significantly.
t-SNE visualization of representation on latent space
Furthermore, Figure 4 illustrates the effect of our method in improving the model’s representation learning on Split-CIFAR100. Specifically, the classes are better clustered, and the separation between them is more distinct. Especially, the classes ’oak tree,’ ’willow tree,’ and ’pine tree’ are divided into clear clusters, rather than being mixed together as in the traditional training strategy, where tasks are trained independently.
The superiority of our proposed method on various types of pretrained backbones.
Table 2 illustrates that our method with the training strategy only consistently outperforms the strongest baseline (HiDE-Prompt) by the gap from about 0.5% to 1.5% in all cases.
Backbone Split CIFAR-100 Split Imagenet-R
TCL HiDE-Prompt TCL HiDE-Prompt
Sup-21K 93.94 92.61 76.41 75.06
iBOT-21K 94.01 93.02 72.12 70.83
iBOT-1K 94.27 93.48 72.80 71.33
DINO-1K 94.12 93.51 69.25 68.11
MoCo-1K 92.32 91.57 64.23 63.77
Table 2: Comparison when using different pretrained backbones
6 Conclusion and Discussion
In this work, we demonstrate the importance of organizing and exploiting data meaningfully rather than lumping it together for training. In particular, we propose organizing data into a tree-like taxonomy based on their label information, helping to identify groups of class data that are likely to confuse models. In this way, we encourage the models to focus and build deeper knowledge about each group, reducing forgetting and motivating more effective learning in subsequent tasks. Additionally, we propose leveraging the initial behavior of pretrained models to unveil a hidden structure of training data, providing an additional perspective to improve performance further. Finally, experimental results demonstrate the effectiveness of these components and our superiority over state-of-the-art baselines.
Limitations.
Despite this novel approach, the quality of the hierarchical taxonomy depends on the quality of expert knowledge. For example, if similar image classes are not assigned to the same leaf group in this label-based taxonomy, the constraint we put on each such group may not perform as expected.
Supplement to “Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning“
Appendix A Experimental Settings
A.1 Datasets
We adopt the following common benchmarks:
  • Split CIFAR-100 [22]: This dataset includes images from 100 different classes, each being relatively small in size. The classes are randomly organized into 10 sequential tasks, with each task containing a unique set of classes.
  • Split ImageNet-R [22]: This dataset contains images from 200 extensive classes. It includes difficult examples from the original ImageNet dataset, as well as newly acquired images that display a variety of styles. The classes are randomly divided into 10 distinct incremental tasks.
  • 5-Datasets [10]: This composite dataset incorporates CIFAR-10 [22], MNIST [23], Fashion-MNIST [52], SVHN [32], and notMNIST [5]. Each of these is treated as a separate incremental task, enabling the evaluation of the impact of substantial variations between tasks.
  • Split CUB-200 [43]: This dataset contains fine-grained images of 200 distinct bird species. It is randomly divided into 10 incremental tasks, each with a unique subset of classes.
A.2 Baselines
In the main paper, we use CL methods with pretrained ViT as the backbone. We group them into (a) the group using a common prompt pool for all tasks, and (b) the group dedicating distinct prompt sets for each task:
(1) L2P [49]: The first prompt-based work for continual learning (CL) suggested using a common prompt pool, selecting the top $k$ most suitable prompts for each sample during training and testing. This approach might facilitate knowledge transfer between tasks but also risks catastrophic forgetting. Unlike our approach, L2P doesn’t focus on training classifiers or setting constraints on features from old and new tasks during training, which may limit the model’s predictability.
(2) DualPrompt [48]: The prompt-based method aims to address L2P’s limitations by attaching complementary prompts to the pretrained backbone, rather than only at input. DualP introduces additional prompt sets for each task to leverage task-specific instructions alongside invariant information from the common pool. However, like L2P, it does not focus on efficiently learning the classification head. Additionally, selecting the wrong prompt ID for task-specific instructions during testing can negatively impact model performance.
(3) OVOR [18]: while using only a common prompt pool for all tasks, this work introduces a regularization method for Class-incremental learning that uses virtual outliers to tighten decision boundaries, reducing confusion between classes from different tasks. Experimental results demonstrate the role of representation learning, which focuses on reducing overlapping between class representations.
(4) S-Prompt++ [47]: S-Prompt was originally proposed for domain-incremental learning, training a separate prompt and classifier head for each task. During evaluation, it infers the domain ID using the nearest centroid from K-Means applied to the training data. To adapt S-Prompt to class-incremental learning (CIL), S-Prompt++ uses a common classifier head for all tasks. However, it shares limitations with DualP, such as efficient learning of the classification head and predicting appropriate prompts during testing.
(5) CODA-Prompt [38]: This prompt-based approach uses task-specific learnable prompts for each task. Similar to L2P, CODA employs a pool of prompts and keys, computing a weighted sum from these prompts to generate the real prompt. The weights are based on the cosine similarity between queries and keys. To avoid task prediction at the end of the task sequence, the weighted sum always considers all prompts. CODA improves over DualP and L2P by optimizing keys and prompts simultaneously, but it still hasn’t addressed the drawbacks mentioned for DualP.
(6) HiDe-Prompt [44]: a recent SOTA prompt-based method that decomposes learning CIL into 3 modules: a task inference, a within-task predictor and a task-adaptive predictor. The second module trains prompts for each task with a contrastive regularization that tries to push features of new tasks away from prototypes of old ones. To predict task identity, it trains a classification head on top of the pretrained ViT. TAP is similar to a fine-tuning step that aims to alleviate classifier bias using the Gaussian distribution of all classes seen so far. However, this method does not declare the relationship between data during training, thereby missing the opportunity to improve model performance.
(7) CPP [25]: This recent SOTA also uses a contrastive constraint to control features of all tasks so far during representation learning and achieves roughly equivalent performance to HiDE on the same settings. Nevertheless, this method still has the advantages that we pointed out in HiDE, which we propose to address in our work.
A.3 Metrics
In our study, we employed two key metrics: the Final Average Accuracy (FAA) and the Final Forgetting Measure (FFM). To define these, we first consider the accuracy on the $i$-th task after the model has been trained up to the $t$-th task, denoted as $A_{i,t}$. The average accuracy of all tasks observed up to the $t$-th task is calculated as $AA_{t}=\frac{1}{t}\sum_{i=1}^{t}A_{i,t}$. Upon the completion of all $T$ tasks, we report the Final Average Accuracy as $\text{FAA}=AA_{T}$. Additionally, we calculate the Final Forgetting Measure, defined as $\text{FFM}=\frac{1}{T-1}\sum_{i=1}^{T-1}\max_{t\in\{1,\ldots,T-1\}}{(A_{i,t}-A_{i,T})}$. The FAA serves as the principal indicator for assessing the ultimate performance in continual learning models, while the FFM evaluates the extent of catastrophic forgetting experienced by the model.
A.4 Implementation Details
Our implementation basically aligns with the methodologies employed in prior research [49, 48, 38]. Specifically, our framework incorporates the use of a pretrained Vision Transformer (ViT-B/16) as the backbone architecture. For the optimization process, we utilized the Adam optimizer, configured with hyper-parameters $\beta_{1}$ set to $0.9$ and $\beta_{2}$ set to $0.999$. The training process was conducted using batches of $24$ samples, and a fixed learning rate of $0.03$ was applied across all models except for CODA-Prompt. For CODA-Prompt, we employed a cosine decaying learning rate strategy, starting at $0.001$. Additionally, a grid search technique was implemented to determine the most appropriate number of epochs for effective training. Regarding the pre-processing of input data, images were resized to a standard dimension of $224\times 224$ pixels and normalized within a range of $[0,1]$ to ensure consistency in input data format. The detailed values of the parameters can be found in our source code.
In Table 1 of the main paper, the results of L2P, DualPrompt, S-Prompt++, CODA-Prompt, and HiDe-Prompt on Split CIFAR-100 and Split ImageNet-R are taken from [44]. Their results on the other two datasets are produced from the official code provided by the authors. For CPP and OVOR, the reported results are also reproduced from their official code. It’s worth noting that the reported forgetting of HiDE is reproduced from their official code.
Appendix B Using ChatGPT to build tree-like taxonomy during a sequence of tasks incrementally
We use the following prompt structure to generate the taxonomies:
Given the label list: [$\cdots$’], provide me the taxonomy from this list, based on their origin, type, and shape, so that the image encoders can recognize their images.
Example output, when the list [$\cdots$’] is ["leopard", "rabbit", "mouse", "camel", "trout", "aquarium_fish", "snake", "rose", "lawn_mower", "bottle"]:
taxonomy = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": ["leopard", "rabbit", "mouse", "camel"]
"Aquatic": ["trout", "aquarium_fish"],
"Reptiles": ["snake"]
"Plants": {
"Flowers": ["rose"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower"],
"Containers": ["bottle"]
Below is an example of generated taxonomies for each task of Split-CIFAR100:
T1 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": ["leopard", "rabbit", "mouse", "camel"]
"Aquatic": ["trout", "aquarium_fish"],
"Reptiles": ["snake"]
"Plants": {
"Flowers": ["rose"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower"],
"Containers": ["bottle"]
T2 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": ["leopard", "rabbit",
"mouse", "camel", "otter"]
"Aquatic": ["trout", "aquarium_fish",
"shark", "seal", "lobster"],
"Reptiles": ["snake"]
"Plants": {
"Flowers": ["rose", "tulip"],
"Trees": ["palm_tree"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower"],
"Containers": ["bottle", "bowl"]
"Vehicles": {
"Wheeled": ["motorcycle"]
"Structures": {
"Buildings": ["skyscraper", "house"]
T3 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": ["leopard", "rabbit", "mouse",
"camel", "otter", "chimpanzee", "squirrel"]
"Aquatic": ["trout", "aquarium_fish", "shark",
"seal", "lobster", "dolphin", "flatfish", "crab"],
"Reptiles": ["snake"]
"Plants": {
"Flowers": ["rose", "tulip"],
"Trees": ["palm_tree", "willow_tree"],
"Fruits": ["sweet_pepper"]
"Environment": {
"Natural Features": ["mountain", "forest"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower"],
"Containers": ["bottle", "bowl"],
"Appliances": ["television"]
"Vehicles": {
"Wheeled": ["motorcycle"]
"Structures": {
"Buildings": ["skyscraper", "house"]
T4 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew"
"Two-legged": ["woman"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab"
"Reptiles": ["snake", "lizard"]
"Plants": {
"Flowers": ["rose", "tulip"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper"]
"Environment": {
"Natural Features": ["mountain", "forest"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl"],
"Appliances": ["television"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"]
T5 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox"
"Two-legged": ["woman"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab"
"Reptiles": ["snake", "lizard"],
"Insects": ["caterpillar", "beetle"]
"Plants": {
"Flowers": ["rose", "tulip"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper"]
"Environment": {
"Natural Features": ["mountain", "forest", "cloud", "plain"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"]
T6 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox", "kangaroo"
"Two-legged": ["woman", "man", "baby"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab"
"Reptiles": ["snake", "lizard"],
"Insects": ["caterpillar", "beetle"],
"Others": ["worm"]
"Plants": {
"Flowers": ["rose", "tulip", "poppy"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper"],
"Fungi": ["mushroom"]
"Environment": {
"Natural Features": ["mountain", "forest", "cloud", "plain"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate", "can"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp", "clock"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle", "pickup_truck"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"],
"Others": ["road"]
T7 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox", "kangaroo", "cattle", "lion"
"Two-legged": ["woman", "man", "baby"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal",
"lobster", "dolphin", "flatfish", "crab", "ray"
"Reptiles": ["snake", "lizard"],
"Insects": ["caterpillar", "beetle", "bee", "cockroach", "spider"],
"Others": ["worm"]
"Plants": {
"Flowers": ["rose", "tulip", "poppy", "sunflower"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper"],
"Fungi": ["mushroom"]
"Environment": {
"Natural Features": ["mountain", "forest", "cloud", "plain"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate", "can"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp", "clock"],
"Furniture": ["bed", "chair"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle", "pickup_truck"],
"Rail": ["train"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"],
"Others": ["road"]
T8 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox", "kangaroo", "cattle", "lion"
"Two-legged": ["woman", "man", "baby"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab", "ray", "whale"
"Reptiles": ["snake", "lizard", "turtle"],
"Insects": ["caterpillar", "beetle", "bee", "cockroach", "spider"],
"Others": ["worm", "snail"]
"Plants": {
"Flowers": ["rose", "tulip", "poppy", "sunflower"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper", "apple", "pear", "orange"],
"Fungi": ["mushroom"]
"Environment": {
"Natural Features": ["mountain", "forest",
"cloud", "plain", "sea"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate", "can"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp", "clock"],
"Furniture": ["bed", "chair", "couch", "table"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle", "pickup_truck", "tractor"],
"Rail": ["train"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"],
"Others": ["road"]
T9 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox", "kangaroo", "cattle",
"lion", "tiger", "wolf", "beaver", "possum", "skunk"
"Two-legged": ["woman", "man", "baby", "boy"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab", "ray", "whale"
"Reptiles": ["snake", "lizard",
"turtle", "crocodile", "dinosaur"],
"Insects": ["caterpillar", "beetle",
"bee", "cockroach", "spider"],
"Others": ["worm", "snail"]
"Plants": {
"Flowers": ["rose", "tulip", "poppy", "sunflower", "orchid"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper", "apple", "pear", "orange"],
"Fungi": ["mushroom"]
"Environment": {
"Natural Features": ["mountain", "forest",
"cloud", "plain", "sea"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate", "can"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp", "clock"],
"Furniture": ["bed", "chair", "couch", "table"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle",
"pickup_truck", "tractor"],
"Air": ["rocket"],
"Rail": ["train"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"],
"Others": ["road"]
T10 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox", "kangaroo", "cattle",
"lion", "tiger", "wolf", "beaver", "possum", "skunk",
"elephant", "bear"
"Two-legged": ["woman", "man",
"baby", "boy", "girl"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab", "ray", "whale"
"Reptiles": ["snake", "lizard",
"turtle", "crocodile", "dinosaur"],
"Insects": ["caterpillar", "beetle",
"bee", "cockroach", "spider", "butterfly"],
"Others": ["worm", "snail"]
"Plants": {
"Flowers": ["rose", "tulip", "poppy", "sunflower", "orchid"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper", "apple", "pear", "orange"],
"Fungi": ["mushroom"]
"Environment": {
"Natural Features": ["mountain", "forest",
"cloud", "plain", "sea"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate", "can", "cup"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp", "clock", "telephone"],
"Furniture": ["bed", "chair", "couch", "table", "wardrobe"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle",
"pickup_truck", "tractor", "bus"],
"Rail": ["train", "streetcar"],
"Air": ["rocket"]
"Structures": {
"Buildings": ["skyscraper", "house", "castle"],
"Bridges": ["bridge"],
"Others": ["road"]
The taxonomies for other datasets are available in our source code.